Nuts to You Shoppe sells cashews for $15 per kg and pecans for $10 per kg. How many kilograms of each should be mixed in order to get 20 kg of a mixture worth $12 per kg?
Please can you show long hand?

Answer by stanbon(75887)

Nuts to You Shoppe sells cashews for $15 per kg and pecans for $10 per kg.
How many kilograms of each should be mixed in order to get 20 kg of a mixture worth $12 per kg?
Please can you show long hand?
-------------
Two variable solution:
Quantity Equation: c + p = 20 kg
Value Equation::: 15c + 10p = 12*20
--------------------------------------------
Multiply thru the Quantity Equation by 15 to get:
15c + 15p = 15*20
---
Subtract that from the Value Equation and solve for "p":
-5p = -3*20
p = 12 kg (amt. of pecans needed in the mixture)
-------------
Since c+p= 20, c = 20-12 = 8 kg (amt. of cashews needed in the mix)
==========================
Cheers,
Stan H.